OpenStudy (impracticaljoker):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). (2 points) y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (impracticaljoker):

@quickstudent @pooja195

OpenStudy (impracticaljoker):

I need help a.s.a.p I need to be done in 25 minutes.

OpenStudy (impracticaljoker):

@AdamK

OpenStudy (adamk):

Last one

OpenStudy (adamk):

Slope is 1-2 / 2 - (-3) = 1/5. Point slope form is y - y1 = m(x-x1).

OpenStudy (impracticaljoker):

thanks

OpenStudy (impracticaljoker):

@AdamK can you help with a couple more?

OpenStudy (blazeryder):

Don't forget to close the question =)